OpenStudy (anonymous):

last 1 I need help with; Solve. x2 - 8x = 20

OpenStudy (anonymous):

the answer would be -2, 10 right?

jhonyy9 (jhonyy9):

x_1,_2=(8+/- sqrt64+80))/2 =(8+/- 12)/2 = -4/2 =-2 and 20/2=10 - so yes you are right